OpenStudy (anonymous):

Trudi and Sean are on a river canoeing. Because of the current of the river, their downstream rate is 250% of their upstream rate. They canoe 3 mi upstream and then return to their starting point. a. Write an expression for the amount of time Trudi and Sean spend canoeing. b. If their upstream rate is 2mi/h, how much time do Trudi and Sean spend canoeing.

OpenStudy (anonymous):

a. Write an expression for the amount of time Trudi and Sean spend canoeing. Let y be the total time taken to canoe and x be the amount of time taken to row upstream. y= x + x/2.5 x is the time taken to row upstream, x/2.5 is the time taken to row downstream. Since downstream will be 250% faster, the time taken will be shorter in that proportion also. b. If their upstream rate is 2mi/h, how much time do Trudi and Sean spend canoeing. Give upstream rate is 2mi/h. Find x, the time taken to row upstream. Since they canoe upstream for 3mi and their rate is 2mi/h Time taken is 3/2 = 1.5 hrs. substitute x = 1.5 to equation of y y= 1.5 + 1.5/2.5 =1.98 hrs

OpenStudy (anonymous):

how did you get 1.98hr?

OpenStudy (anonymous):

y= 1.5 + 1.5/2.5 =1.98 hrs

OpenStudy (anonymous):

isn't 1.5 divided by 2.5= 0.6?

OpenStudy (anonymous):

Hmm i don't know how i made that error, punch calculator wrongly. Yep you are right!